FAX Cover Page

FAX Page Count: 2

17 March 2000, 9:42 A.M.

To:

Pyotr Ilyitch Tchyfontski Research Associate Professor Division of Fontal Conversion St. Leninberg University of the New State Telephone: Telephone: Telephone: Telephone: FAX:

From:

# 5 Fifth Ave.

New Font City, New Fontiac

United States of Fontagraphy 12345-6789

Ludwig von Beetfonten # 5 Fifth Ave. New Font City, New Fontiac United States of Fontagraphy 12345-6789 Telephone: 030-1234567 Telephone: Telephone: Telephone: FAX: 030-55535552 Pyotr Ilyitch Tchyfontski Research Associate Professor Division of Fontal Conversion St. Leninberg University of the New State

Tovarich Pyotr:

A line that will differ from case to case.In terms of the relationships between font descriptions and musical typography, it seems to me that only in the case of the early 1800s can even the slightest case be made for a relationship.

Yours in fontal health,

Ludwig von Beetfonten

Research Associate Professor Division of Font Music Insertions Vienna University

Your font is my font!!

FAX Cover Page

FAX Page Count: 2

17 March 2000, 9:42 A.M.

To:

Carmen “Cacciatore” Rodriguez Full Assistant Associate Professor Division of Chicken Informatics Cuba National University Guantanemo Bay, Cuba Telephone: Telephone: Telephone: Telephone: FAX:

From:

Ludwig von Beetfonten # 5 Fifth Ave. New Font City, New Fontiac United States of Fontagraphy 12345-6789 Telephone: 030-1234567 Telephone: Telephone: Telephone: FAX: 030-55535552 Carmen “Cacciatore” Rodriguez Full Assistant Associate Professor Division of Chicken Informatics Cuba National University Guantanemo Bay, Cuba

Senora Rodriguez:

Of course, it doesn’t have to, but it will anyway.In terms of the relationships between font descriptions and musical typography, it seems to me that only in the case of the early 1800s can even the slightest case be made for a relationship.

Yours in fontal health,

Ludwig von Beetfonten

Research Associate Professor Division of Font Music Insertions Vienna University

Your font is my font!!